OpenStudy (anonymous):

How do the graphs of y= 1/x and y= 5/x+6? PLEASE EXPLAIN.

OpenStudy (anonymous):

I really need some help with this question

OpenStudy (anonymous):

-_- i kinda getting pissed off here!

OpenStudy (anonymous):

the question is incomplete

OpenStudy (anonymous):

it is not

OpenStudy (anonymous):

how do you draw the graphs? because "how do the graphs of ..." isnt asking for anything

OpenStudy (anonymous):

i tried doing this problem on my calculator and im still confused

OpenStudy (anonymous):

-_-

OpenStudy (anonymous):

are you trying to graph it on a calculator? or are you trying to find the intersection point of the 2 graphs?

OpenStudy (anonymous):

im trying to graph it

OpenStudy (anonymous):

ok then is your problem, that there is no x=0?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok if x= 1/2, y=? if x=1/4, y=? if x= 1/8, y=? as we get closer and closer to zero, you will notice that y will become a really large number in terms of limits, as you approach x= 0 from the right, y approaches infinity unfortunately, your calculator is unable to handle the concept of infinity or really really really large numbers which is why you will get an error when you try to divide by zero or a really really really small fraction

OpenStudy (anonymous):

oh...

OpenStudy (anonymous):

now try to do the same for negative numbers -1/2, -1/4, -1/8

OpenStudy (anonymous):

as the number becomes an even smaller negative fraction, you will notice the result "y" becomes a higher and higher negative number as x approaches 0 from the left, y approaches negative infinity

OpenStudy (anonymous):

im still confused

OpenStudy (anonymous):

nvm i got help

OpenStudy (anonymous):

just looking at y=1/x if x=100, y=1/100 x=10, y=1/10 x=1 , y=1 x=.1 , y=10 x=.01,y=100 etc etc etc if we were to graph these points we would get something that looks like